OpenStudy (anonymous):

multple choice: which value of r gives the line passing through (3,2) and (r,-4) a slope of 3/2? A. -6 B. -1 C. 7 D. 12

OpenStudy (callisto):

\[slope =\frac{y_2-y_1}{x_2-x_1}= \frac{2-(-4)}{3-r} = \frac{3}{2}\]\[\frac{2+4}{3-r} = \frac{3}{2}\] Can you continue?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you

OpenStudy (callisto):

welcome!~